A typical wall outlet voltage is 120 volts in the United States. Personal MP3 players require much smaller voltages, typically 230.5 mV. If the number of turns in the primary coil is 4686, calculate the number of turns on the secondary coil of the adapter transformer.
Answer #1

N=4686*230.5/(120*1000)

=9
Answer #2

v1/v2= n1/n2


n2= v2*n1/v1

= 0.2305*4686/120
= 9.001volts

=9volts